# MacRAG: Compress, Slice, and Scale-up for Multi-Scale Adaptive Context RAG

### v1

@@ 3.2.5 Neighbor Chunk Propagation & Merging
Within each of the top-$k_{2}$ distinct documents, we focus on chunks $c_{i}$ both in $\mathcal{D}_{q,k_{2}}$ and $\mathcal{C}^{{}^{\prime}}_{q,k_{2}}$, as these represent the most relevant content segments. For such a chunk $c_{i}$, we define $\mathcal{N}_{h}(c_{i})$ which consists of $c_{i}$’s neighbors within $h$-hops and $c_{i}$ itself, and process extension of $c_{i}$ to get $\mathcal{N}(c_{i})$ by using its associated indices,
$\mathcal{N}_{h}(c_{i})=\;\bigl{\{}\,c_{i-h},\dots,c_{i},\dots,c_{i+h}\bigr{\}}.$ (7)
We then merge $\mathcal{N}_{h}(c_{i})$ within the same ranked document to form the final merged chunk for adaptive long context to the query $q$:
$c^{merge}_{q,d}~{}=~{}$ $\mathrm{Merge}\bigl{(}\mathcal{N}_{h}(c_{i})\bigr{)},$ (8)
$\text{where}~{}d\in$ $\mathcal{D}_{q,k_{2}},~{}c_{i}\in d~{}and~{}c_{i}\in\mathcal{C}^{{}^{\prime}}_{q,k_{2}}.$
As described in Equation 8, from the top-$k_{2}$ documents in $\mathcal{D}_{q,k_{2}}$, we can select top-$k_{2}$ merged chunks $c^{merge}_{q,d}$ by considering its order along with top document order with $r_{q,d}$. Then, finally we have top-$k_{2}$ merged chunks as
$\mathcal{C}_{\mathrm{final}}~{}=~{}\mathrm{TopK}_{k_{2}}\!\Bigl{(}c^{merge}_{q,d}\Bigr{)}.$ (9)
The final merged chunks $\mathcal{C}_{\text{final}}$ provide diverse views across the top-$k_{2}$ documents while optimizing context length (Zhao et al., 2024), combining the most relevant content with minimal redundancy.
Mode Description
R&L Single-step generation using complete documents from top-$k_{2}$ chunks. Prone to “Lost in the Middle” effect and inefficient for long documents.
Full_E&F Multi-step generation combining Full_Ext’s extracted content and Fil’s filtered chunks. Balances coverage and precision, but expensive for long doc.
R&B_E&F Multi-step generation combining R&B_Ext’s extracted content and Fil’s filtered chunks. Optimizes precision-recall trade-off w/o processing full doc.
@@ 3.3 Single/Multi-step Post Retrieval Generation
Our retrieval framework is modular, enabling integration with various post-retrieval generation approaches. These include both vanilla single-step generation directly using the retrieved top-$k_{2}$ chunks in a single pass and iterative multi-step generation as in LongRAG (Zhao et al., 2024), which introduces CoT-guided filter (which checks chunk-level relevance) and LLM-augmented extractor (which locates global context in original passages). Following the comprehensive exploration of single and multi-step generation variants in LongRAG, we evaluate MacRAG’s performance when combined with different post-retrieval generation methods. Specifically, we investigate seven generation methods, as detailed in Table 1. Five of these methods—R&B, R&L, Full_Ext, Fil, and Full_E&F—are adopted directly from LongRAG. Among these, R&L, Full_Ext, and Full_E&F process the entire set of top-$k_{2}$ documents, which can introduce computational overhead when handling very long texts. To address this limitation while maintaining an optimal balance between precision and recall, we introduce two new variants: R&B_Ext and R&B_E&F, to reduce this overhead while maintaining precision-recall balance.
@@ 3.4 Implementation Complexity and System Efficiency
MacRAG enhances simple overlapping-chunk indexing by first chunking each document with partial overlaps, then compressing each chunk through summarization, and finally slicing it. These processes are performed once during offline preparation, eliminating query-time overhead. Moreover, this document-level independence simplifies updates and streamlines database operations, avoiding the complexities associated with hierarchical multi-clustering Sarthi et al. (2024); Zhang et al. (2025).
At inference time, MacRAG constructs a real-time effective long context by mapping slices back to their parent chunks and documents using lightweight index-based lookups rather than costly tree traversals. This design leverages metadata dictionaries (document ID, chunk ID, token offsets) to enable modular neighbor merging with minimal overhead. A bounded context size is enforced through parameters $(k_{2}\times\alpha)$ and chunk dimensions, ensuring scalability to enterprise-level corpora without degrading performance. This controlled scalability, along with real-time context assembly, makes MacRAG highly valuable for massive long-document collections, “almost infinite context” in advanced personalization, and iterative and agentic RAGs.
@@
Table 3: Consolidated results for four performance metrics across three datasets, comparing LongRAG and MacRAG using the Gemini-1.5-pro and GPT-4o models. The results are averaged over seven generation settings in Table 1.
@@ 4.1 Experimental Setup
Benchmark Datasets and Challenges:
We evaluate MacRAG on three multi-hop question answering benchmarks from the expanded LongBench framework (Bai et al., 2023): HotpotQA, 2WikimultihopQA, and Musique. These datasets have become the standard evaluation suite for hierarchical retrieval research, including recent systems such as SIRERAG Zhang et al. (2025), RAPTOR Sarthi et al. (2024), and LongRAG Zhao et al. (2024). Their LongBench variants specifically intensify the “Lost in the Middle" phenomenon by introducing near-duplicate and tangential passages, creating challenging retrieval environments where critical connecting information becomes obscured by surrounding context. Each dataset introduces distinct multi-hop reasoning challenges that systematically test different aspects of retrieval capability: HotpotQA features bridging and comparison queries requiring evidence synthesis across multiple paragraphs; 2WikiMultihopQA expands partial excerpts to complete articles, significantly increasing the risk that essential connecting information becomes buried within expansive context; and Musique embeds nested sub-questions within tangential sections, challenging systems to identify and connect scattered bridging facts across document boundaries. These characteristics directly test a retrieval system’s ability to handle fragmented evidence, reconcile overlapping information, and identify critical bridging facts despite contextual noise (Yue et al., 2024; Leng et al., 2024), precisely the challenges that MacRAG’s multi-scale architecture addresses through hierarchical indexing and adaptive retrieval.
Models and Evaluation Metrics:
Based on SIRERAG’s comparative analysis (Table 4 in Zhang et al. (2025)), we focus our evaluation on RAPTOR Sarthi et al. (2024) and LongRAG Zhao et al. (2024) as representative state-of-the-art hierarchical RAG systems. SIRERAG’s analysis establishes that GraphRAG underperforms on these specific multi-hop tasks, while RAPTOR, HippoRAG, and SIRERAG achieve comparable performance in terms of F1-score on multiple multi-hop question answering datasets Zhang et al. (2025). Among these top-performing systems, we selected RAPTOR and LongRAG due to their better balance between performance and computational efficiency, offering lower computational costs and easier maintenance compared to SIRERAG’s higher overhead.
We conducted experiments using three prominent long-context LLMs: two widely adopted proprietary models (GPT-4o and Gemini-1.5-pro) and one popular open-source model (Llama-3.1-8B-instruct). This diverse selection enables assessment of MacRAG’s generalizability across different model architectures and parameter scales.
We measure performance through four complementary metrics: Exact Match, F1-score, Precision, and Recall, providing a multidimensional view of answer quality. We additionally track average context length to quantify MacRAG’s efficiency in reducing input redundancy while maintaining or improving information coverage.
To ensure fair comparison across all methods (MacRAG, LongRAG, and RAPTOR), we basically applied the same re-ranking model (marco-miniLM) for all RAG methods.
Comparative Framework and Hyperparameters:
To ensure rigorous comparative analysis across all methods (MacRAG, LongRAG, and RAPTOR), we used the same re-ranking model (marco-miniLM) for all RAG methods. We also adopted the optimal hyperparameter configuration from LongRAG Zhao et al. (2024), and used identical hyperparameters across all methods: $k_{1}=100$ chunks retrieved initially, $k_{2}=7$ chunks selected after reranking, except in ablation studies. We maintain identical prompt templates across all methods to eliminate prompt engineering as a confounding variable.
@@
For MacRAG-specific parameters, we varied the scale-up factor $\alpha\in\{1,4\}$ and the neighbor hop count $h\in\{0,1\}$, except during ablation studies. The ablation study and performance analysis in Section 4.2 demonstrate both robustness and the soundness of the architecture design of MacRAG, enhancing multi-hop reasoning capability.
@@
This carefully controlled experimental framework ensures that observed performance differences can be attributed specifically to MacRAG’s architectural innovations, particularly hierarchical document representation and adaptive, multi-scale retrieval rather than variations in hyperparameter optimization or underlying retrieval technologies. By maintaining consistency while isolating architectural factors, we can precisely measure the impact of multi-scale retrieval on multi-hop reasoning performance.

### v2

@@ 3.2.5 Neighbor Chunk Propagation & Merge for Input Context Construction
Within each of the top-$k_{2}$ distinct documents, we focus on chunks $c_{i}$ both in $\mathcal{D}_{q,k_{2}}$ and $\mathcal{C}^{{}^{\prime}}_{q,k_{2}}$, as these represent the most relevant content segments. For such a chunk $c_{i}$, we define $\mathcal{N}_{h}(c_{i})$ which consists of $c_{i}$’s neighbors within $h$-hops and $c_{i}$ itself, and process extension of $c_{i}$ to get $\mathcal{N}(c_{i})$ by using its associated indices, $\mathcal{N}_{h}(c_{i})=\;\bigl{\{}\,c_{i-h},\dots,c_{i},\dots,c_{i+h}\bigr{\}}$. We then merge $\mathcal{N}_{h}(c_{i})$ within the same ranked document to form the final merged chunk for adaptive long context to the query $q$: $c^{merge}_{q,d}=\mathrm{Merge}\bigl{(}\mathcal{N}_{h}(c_{i})\bigr{)}$, where $d\in\mathcal{D}_{q,k_{2}}$, $c_{i}\in d$, and $c_{i}\in\mathcal{C}^{{}^{\prime}}_{q,k_{2}}$. From the top-$k_{2}$ documents in $\mathcal{D}_{q,k_{2}}$, we can select top-$k_{2}$ merged chunks $c^{merge}_{q,d}$ by considering their order along with the top document order with $r_{q,d}$. Then, finally, we have top-$k_{2}$ merged chunks as $\mathcal{C}_{\mathrm{final}}=\mathrm{TopK}_{k_{2}}\!\Bigl{(}c^{merge}_{q,d}\Bigr{)}$. The final merged chunks $\mathcal{C}_{\text{final}}$ provide diverse views across the top-$k_{2}$ documents while optimizing context length (Zhao et al., 2024), combining the most relevant content with minimal redundancy.
@@ 3.3 Single/Multi-step Post Retrieval Generation
Our retrieval framework is modular, enabling integration with various post-retrieval generation approaches. These include both vanilla single-step generation directly using the retrieved top-$k_{2}$ chunks in a single pass and iterative multi-step generation as in LongRAG (Zhao et al., 2024), which introduces a CoT-guided filter (which checks chunk-level relevance) and an LLM-augmented extractor (which locates global context in original passages). Following the comprehensive exploration of single and multi-step generation variants in LongRAG, we evaluate MacRAG’s performance when combined with different post-retrieval generation methods. Specifically, we investigate seven generation methods, as detailed in Table 4. Five of these methods (R&B, R&L, Full_Ext, Fil, and Full_E&F) are adopted directly from LongRAG. Among these, R&L, Full_Ext, and Full_E&F process the entire set of top-$k_{2}$ documents, which can introduce computational overhead when handling very long texts. To address this limitation while maintaining an optimal balance between precision and recall, we introduce two new variants: R&B_Ext and R&B_E&F, to reduce this overhead while maintaining precision-recall balance.
@@ 3.4 Implementation Complexity and System Efficiency
MacRAG enhances simple overlapping-chunk indexing by first chunking each document with partial overlaps, then compressing each chunk through summarization, and finally slicing it. These processes are performed once during offline preparation, eliminating query-time overhead. Moreover, this document-level independence simplifies updates and streamlines database operations, avoiding the complexity associated with hierarchical multi-clustering Sarthi et al. (2024); Zhang et al. (2025).
At inference time, MacRAG constructs a real-time effective long context by mapping slices back to their parent chunks and documents using lightweight index-based lookups rather than costly tree traversals. This design leverages metadata dictionaries (document ID, chunk ID, token offsets) to enable modular neighbor merging with minimal overhead. A bounded context size is enforced through parameters $(k_{2}\times\alpha)$ and chunk dimensions, ensuring scalability to enterprise-level corpora without degrading performance. This controlled scalability and real-time context assembly underscore MacRAG’s value for massive long-document collections, “almost infinite context" scenarios in advanced personalization, and as a robust foundation for iterative RAG or agentic systems.
@@
Table 2: Consolidated results for four metrics across three datasets, comparing LongRAG and MacRAG using the Gemini-1.5-pro and GPT-4o models. The results are averaged over seven generation settings in Table 4.
@@ 4.1 Experimental Setup
We evaluate MacRAG on three challenging multi-hop question answering benchmarks from LongBench (Bai et al., 2023): HotpotQA, 2WikiMultihopQA, and Musique. These datasets, with their tangential passages and obscured connecting information, specifically test a system’s ability to handle the "Lost in the Middle" phenomenon and perform robust multi-hop reasoning—precisely the challenges MacRAG addresses. Section A.2 in the Appendix provides a detailed explanation about these datasets.
Models and Evaluations.
Our evaluations employ three prominent LC LLMs: GPT-4o (OpenAI, 2024), Gemini-1.5-pro (Team et al., 2024), and the open-source Llama-3.1-8B-instruct (Dubey et al., 2024a). Performance is measured using Exact Match (EM), F1-score, Precision, and Recall. We primarily compare MacRAG against strong baselines RAPTOR (Sarthi et al., 2024) and LongRAG (Zhao et al., 2024), selected for their representative state-of-the-art performance and efficiency balance in hierarchical RAG (as discussed in Section 2). To ensure rigorous comparative analysis across all RAG methods (MacRAG, LongRAG, and RAPTOR) use the same reranker (marco-miniLM), and LongRAG’s reported optimal hyperparameters ($k_{1}=100$ initial chunks, $k_{2}=7$ final chunks, unless in ablation). We also applied seven different generation methods described in Appendix A.1, and Table 4 for detailed comparisons.
@@
MacRAG-specific parameters are scale-up factor $\alpha\in\{1,4\}$ and the neighbor hop count $h\in\{0,1\}$, except during ablation studies. The ablation study and performance analysis in Section 4.2 demonstrate both robustness and the soundness of the architecture design of MacRAG, enhancing multi-hop reasoning capability.
@@
This controlled setup ensures that performance differences result from MacRAG’s architectural innovations, particularly its hierarchical representation and adaptive multi-scale retrieval, rather than from variations in hyperparameters or retrieval components. By isolating architectural factors, we can clearly assess the impact of multi-scale retrieval on multi-hop reasoning.
@@ 4.2 Main Results
HotpotQA
LongRAG 65.46 69.40 66.60 63.32 66.20 66.08 66.09 66.16
2WikimultihopQA
LongRAG 59.97 62.37 65.35 60.68 65.89 62.08 62.47 62.69
Musique
LongRAG 38.98 41.90 43.64 37.72 43.83 42.00 41.97 41.43
Table 3: Ablation study of MacRAG with GPT-4o and F1-score, using the same $k_{1}$, $k_{2}$, and reranker.
Comparison with RAPTOR.
While RAPTOR Sarthi et al. (2024), a summarization and hierarchical clustering-based approach, shows gains on standard datasets, it struggles with retrieval quality on LongBench versions Bai et al. (2023) as shown in Table 1. Specifically, RAPTOR with Llama-3.1-8B shows performance decreases in 2WikimultihopQA and Musique, potentially associated with retrieval difficulty on Musique-Normal in (Yue et al., 2024). These results highlight the importance of both precise retrieval and effective long-context construction for achieving consistent gains across various multi-hop reasoning and long-context datasets.
Integration with Multi-step Generation Method.
Integrating MacRAG with the LongRAG (E&F) generation method across multiple LLMs (Llama-3.1-8B, Gemini-1.5-pro, and GPT-4o) yields substantial performance improvements, as shown in Table 1. This underscores MacRAG’s versatility and robustness as a retrieval approach when paired with the LongRAG generation method. By constructing an effective long context, MacRAG demonstrates advantages in multi-step generation with LongRAG (E&F), producing consistent and meaningful gains across three datasets and three LLMs. In contrast, it is not straightforward to extend RAPTOR in combination with LongRAG (E&F). As further illustrated in Table 1, the performance gains for Gemini-1.5-pro and GPT-4o indicate that pairing MacRAG with more powerful LLMs can achieve even greater improvements, despite already high F1-scores obtained by other frameworks under these settings. We further observe that summarization during indexing improves retrieval quality. Our preliminary experiments show a 7.2% gain in precision when using summaries instead of raw slices.
Robustness across Generation Methods.
In addition to MacRAG’s robust gains in both vanilla single-step (R&B) and multi-step generation (E&F), MacRAG consistently shows strong advantages across the four metrics listed in Table 2, covering all seven generation modes in Table 4. These substantial improvements across various datasets, metrics, and LLMs demonstrate the effectiveness of MacRAG’s long-context retrieval in enhancing both precision and recall for RAG. Notably, MacRAG achieves significant gains on the Musique datasets, which Yue et al. (2024) identifies as having challenging retrieval conditions. Beyond the average gains in these four metrics, Figure 2 further confirms MacRAG’s robust benefits in all seven single-/multi-step generation schemes in Table 4 in Appendix A.1. Finally, regardless of the choice of $k_{2}$ or reranker, MacRAG maintains substantial advantages in all test settings in Table 5.
Enhanced Gains with Stronger Models.
When applied to the more powerful Gemini-1.5-pro and GPT-4o models, MacRAG showcases its capacity to further improve performance by optimizing the retrieval process and building more relevant, informative contexts. Its use of hierarchical chunking and slicing, coupled with adaptive propagation and merging, maintains high precision while expanding coverage. This leads to more efficient handling of long documents, reduced computational overhead, and higher-quality generated answers overall.
Ablation Study.
Table 3 presents an ablation study demonstrating the effectiveness of MacRAG’s key components: Propagation and Merging, Scaling-up, and Hierarchical Slicing Retrieval. Each component contributes to cumulative performance gains. The combination of propagation and merging with the scaling-up mechanism increases the coverage of relevant contexts for multi-hop reasoning tasks. Figure 3 further illustrates the benefits of scaling up promising candidates with $\alpha=[2,3,4]$, showing how this approach effectively balances context lengths while incorporating additional relevant candidates. We observe that performance gains remain consistent across different $k_{2}$ values and rerankers, indicating robustness. Although the parameters $(k_{1},k_{2},\alpha,h)$ are fixed, MacRAG adaptively merges nearby slices depending on partial relevance signals. Removing this adaptivity leads to a 5% F1 drop, as shown in our ablations.
4.3 Generation Schemes and Input Lengths
Figure 4: Cumulative LLM’s input context lengths for single/multi-step generation methods in Table 4.
While multi-step generation methods (e.g., LongRAG’s Full_E&F) improve answer quality, they often increase cumulative input length. Figure 4 compares cumulative input lengths for single and multi-step generation methods in Table 4, given single retrieval with specific settings, highlighting that R&L, Full_Ext, and Full_E&F accumulate the most overhead by relying on entire documents. In contrast, MacRAG’s final $k_{2}$ chunks, though potentially larger than basic R&B chunks, preserve bounded context well below full-document with long lengths.
As Figure 2 and Table 3 illustrate, MacRAG consistently outperforms baselines under both R&B_Ext and R&B_E&F settings. Due to MacRAG’s advantage in effective and stable long context construction, even our intermediate R&B_E&F variant demonstrates significant improvements as shown in Table 3. For example, on 2WikiMultihopQA (GPT-4o), MacRAG (R&B_E&F) achieves an F1 of 72.43 versus LongRAG (Full_E&F) at 65.89 (+9.93% relative gain), while using 8.19% less context. On Musique, the intermediate variant of MacRAG (R&B_E&F) reaches 45.84 compared to LongRAG (Full_E&F)’s 43.83 (+4.58% relative) with a 45.87% reduction in context length. This highlights MacRAG’s efficiency in evidence gathering and context construction.
4.4 Efficiency and Latency Analysis
MacRAG maintains sub-second efficiency despite constructing richer multi-scale contexts. On average, retrieval and reranking take 0.23 seconds per query, which is 38% faster than RAPTOR’s 0.37 seconds. With Llama 3.1 8B, total inference latency is 0.99 seconds (0.23s retrieval and 0.76s generation), compared to 0.80 seconds for RAPTOR (0.37s and 0.43s, respectively). This slight increase in generation time is offset by improved retrieval precision and higher answer quality on R&B queries.
This efficiency is enabled by MacRAG’s lightweight, index-based context construction and shared reranking modules, which preserve the same level of computational complexity as baseline retrieval pipelines. MacRAG introduces a one-time offline summarization step during indexing, similar to RAPTOR and SIRERAG. This step runs in parallel with standard RAG indexing, adding no runtime latency and supporting efficient index maintenance. As shown in Table 3, Figure 4, and Subsection 4.3, MacRAG achieves a better trade-off than LongRAG between input context length, answer quality, and overall computational cost.
5 Conclusion
In this work, we introduced Multi-scale Adaptive Context RAG (MacRAG), a novel framework designed to address key trade-offs in RAG involving precision, coverage, and computational efficiency for long-context question answering and multi-hop reasoning. By dynamically constructing query-adaptive long contexts through real-time, multi-scale retrieval, MacRAG demonstrably improves both precision and recall while maintaining operational efficiency. Our extensive experiments validated MacRAG’s advantages across diverse datasets and both single- and multi-step generation paradigms, showcasing consistent and notable performance gains. These findings establish MacRAG as a potent and efficient core module for advancing next-generation iterative and agentic RAG systems, with significant potential for robust enterprise-level applications.
Limitations
While MacRAG demonstrates strong performance with its modular and efficient design, several avenues offer potential for further extension. The current offline summarization and indexing strategy, though effective and its costs amortized, could benefit from adaptive or online variants to better handle rapidly evolving corpora. Furthermore, while the use of fixed hyperparameters such as chunk size and expansion scale has shown robustness, dynamically adjusting these parameters based on query complexity presents a promising research direction. Finally, tighter integration of MacRAG with sophisticated agentic or multi-round retrieval workflows could further enhance its applicability in highly interactive and real-time RAG scenarios.
Appendix A Appendix
A.1 Generation Modes
Table 4 summarizes different generation modes. In addition to MacRAG’s robust gains in both vanilla single-step (R&B) and multi-step generation (E&F), MacRAG consistently shows strong advantages across the four metrics listed in Table 2, covering all seven generation modes in Table 4. These substantial improvements across various datasets, metrics, and LLMs demonstrate the effectiveness of MacRAG’s long-context retrieval in enhancing both precision and recall for RAG. Notably, MacRAG achieves significant gains on the Musique datasets, which Yue et al. (2024) identifies as having challenging retrieval conditions. Beyond the average gains in these four metrics, Figure 2 further confirms MacRAG’s robust benefits in all seven single-/multi-step generation schemes in Table 4 in Appendix A.1. Finally, regardless of the choice of $k_{2}$ or reranker, MacRAG maintains substantial advantages in all test settings in Table 5.
Mode Description
R&L Single-step generation using complete documents from top-$k_{2}$ chunks. Prone to the “Lost in the Middle” effect and inefficient for long documents.
Full_E&F Multi-step generation combining Full_Ext’s extracted content and Fil’s filtered chunks. Balances coverage and precision, but is expensive for long documents.
R&B_E&F Multi-step generation combining R&B_Ext’s extracted content and Fil’s filtered chunks. Optimizes precision-recall trade-off w/o processing the full document.
A.2 Benchmark Datasets and Challenges
The LongBench variants of HotpotQA, 2WikimultihopQA, and Musique datasets have become the standard evaluation suite for hierarchical retrieval research, including recent systems such as SIRERAG Zhang et al. (2025), RAPTOR Sarthi et al. (2024), and LongRAG Zhao et al. (2024). They specifically intensify the “Lost in the Middle" phenomenon by introducing near-duplicate and tangential passages, creating challenging retrieval environments where critical connecting information becomes obscured by surrounding context. Each dataset introduces distinct multi-hop reasoning challenges that systematically test different aspects of retrieval capability: HotpotQA features bridging and comparison queries requiring evidence synthesis across multiple paragraphs. 2WikiMultihopQA expands partial excerpts to complete articles, significantly increasing the risk that essential connecting information becomes buried within expansive context. Musique embeds nested sub-questions within tangential sections, challenging systems to identify and connect scattered bridging facts across document boundaries. These characteristics directly test a retrieval system’s ability to handle fragmented evidence, reconcile overlapping information, and identify critical bridging facts despite contextual noise (Yue et al., 2024; Leng et al., 2024), precisely the challenges that MacRAG’s multi-scale architecture addresses through hierarchical indexing and adaptive retrieval.
A.3 Target Baseline Methods
Based on SIRERAG’s comparative analysis (Table 4 in Zhang et al. (2025)), we focus our evaluation on RAPTOR Sarthi et al. (2024) and LongRAG Zhao et al. (2024) as representative state-of-the-art hierarchical RAG systems. SIRERAG’s analysis establishes that GraphRAG underperforms on these specific multi-hop tasks, while RAPTOR, HippoRAG, and SIRERAG achieve comparable performance in terms of F1-score on multiple multi-hop question answering datasets Zhang et al. (2025). Among these top-performing systems, we selected RAPTOR and LongRAG because they balance performance and computational efficiency better, offering lower computational costs and easier maintenance than SIRERAG’s higher overhead.
A.4 Adaptive Retrieval Mechanism Analysis
Query-dependent Adaptive Context Expansion:
A.5 Preprocessing Efficiency of MacRAG
Chunk Size and Token Budget:
MacRAG uses original chunk sizes of approximately 400 tokens (around 1500 characters), which fall within the 200–500 token range used in prior work such as LongRAG even for re-ranking on retrieved chunks. The average sizes of summarized chunks are around 800-1000 characters. We applied slices into summarized chunks for initial retrieval to improve precision further with two slice sizes around 450 characters with 300 character overlaps or 600 characters with 450 character overlaps. These sizes allow MacRAG to maintain sub-second retrieval and reranking times while keeping the context structure manageable for downstream language models.
Offline Summarization Cost:
A.6 Latency Comparison between MacRAG and RAPTOR
Retrieval and Reranking Latency:
Generation Time and Trade-offs:
Total Latency Comparison:
A.7 Efficiency Comparison between MacRAG and LongRAG
A.8 Discussion on Effectiveness of MacRAG’s Retrieval and Constructed Long Context
A.9 Discussion on Single, Multi-Step, Iterative, and Agentic Generation with MacRAG
A.10 Discussion on Graph-Enhanced MacRAG